OpenStudy (anonymous):

If sin Θ = negative square root 3 over 2 and π < Θ < 3 pi over 2, what are the values of cos Θ and tan Θ?

OpenStudy (anonymous):

\[\sin \theta= -\sqrt{3/2} \]

OpenStudy (anonymous):

The rest looks fine

OpenStudy (phi):

are you sure it's not \[ \sin\theta = - \frac{\sqrt{3}}{2} \]?

OpenStudy (anonymous):

*facepalm* It's so sad that you're right, and you don't even get to see the problem. But yes, you are correct.

OpenStudy (phi):

pi to 3pi/2 is the 3rd quadrant draw a triangle in the 3rd quadrant, label the angle theta and label the sides of the triangle knowing sin x = opp/hyp

OpenStudy (phi):

it should look like this